# Inspection-Guided Randomization: A Flexible and Transparent Restricted Randomization Framework for Better Experimental Design

### v1

@@ 5 Discussion
We introduced Inspection-Guided Randomization as a flexible framework for performing restricted randomization with diverse design desiderata and used illustrative examples to showcase its usefulness in different experimental settings. The core motivation for IGR, as for any restricted randomization, is to prevent the possibility of choosing a bad randomization allocation that leads to an effect estimate that deviates dramatically from the true value of the estimand. For obviously bad allocations, there is little reason not to discard them; this has been well-understood since the early days of the experimental design literature. In a conversation between Cochran and Fisher (recounted by Rubin and Cochran), Fisher is reported to have said that, confronted with a randomized allocation that “exhibited substantial imbalance on a prognostically important baseline covariate,” he would “of course” rerandomize if the experiment had not started yet (Rubin,, 2008).
@@
Historically, the focus of restricted randomization has been on achieving covariate balance in two-arm experiments that compare a single treatment intervention to a control or status quo. However, some interventions, often ones that are behavioral, need to be evaluated through more complicated restricted designs. Researchers across several disciplines, including medicine, education, psychology, and computer science, have recognized this need and have independently applied principles of the IGR framework to designing experiments, albeit without the formal structure of IGR (Murnane et al.,, 2023; Cho et al.,, 2021, 2022; Kizilcec et al.,, 2022; Zahrt et al.,, 2023). This suggests that IGR is conceptually straightforward and intuitive, and it also points to how an explicit IGR framework can help to bring consistency to existing efforts to expand the scope of restricted randomization.
In this work, we showed how the IGR framework can be used to design experiments in two settings where existing restricted randomization frameworks are not immediately applicable. First, in Vignette 1 (Section 4.1), we applied IGR to the design of group formation experiments, where effect sizes may exhibit dependence on group composition. Recent work has highlighted that peer context should be taken into consideration as a driver of treatment effect heterogeneity for behavioral and psychological interventions (Yeager et al.,, 2019; Walton and Yeager,, 2020; Kizilcec et al.,, 2020; Kizilcec and Cohen,, 2017). Understanding how an intervention depends on peer context has important implications for understanding causal mechanisms, for devising policies that target certain contexts, and for designing concurrent interventions that can be used to manipulate contexts. IGR can also be broadly compatible with definitions of peer contexts beyond the one used in Vignette 1, such as those that involve more than one attribute or those that are continuous functions of peer characteristics.
@@
In Vignette 2 (Section 4.2), we applied IGR to the design of experiments with interference. Experimental designs that seek to mitigate interference often do so through a form of cluster randomization, either over naturally existent clusters or over algorithmically detected ones (Ugander et al.,, 2013; Ugander and Yin,, 2023). Without IGR, the study analyst chooses between randomizing at the more aggregated level to limit interference, thus sacrificing power and covariate balance, or randomizing at the more granular level to increase efficiency, thus permitting more interference bias. With IGR, the analyst can construct fitness functions that combine both interference and balance metrics, thus reducing interference bias while maintaining efficiency. IGR can also be used even if we have limited knowledge of interference structure and need to rely on an approximate measure of interference. While we explored a specific setting where interference occurred via neighborhood exposure in a network of interactions, IGR can easily be applied to other interference structures by modifying the inspection metric used to measure the extent of anticipated interference.
Fundamentally, IGR compels the study analyst to invest time and care in the design phase of the experiment. The Specification and Evaluation & Adaptation steps in particular require considerable thoughtfulness, from specifying the study’s design desiderata, to developing the inspection metrics used to measure desiderata satisfaction, to constructing and evaluating the fitness function and restriction rule that characterize the restricted assignment mechanism. Attentiveness to design can produce large payoffs in the bias and precision of effect estimates. While using a tailored estimator in the analysis phase can sometimes produce similar payoffs, focusing efforts on the design phase has additional merits. For one, even the most clever estimators may not be able to salvage a study that was performed under an extremely poor choice of randomization allocation. In addition, as previously mentioned, the Pre-registration step precludes statistical hacking, as long as the analyst is held accountable to doing randomization inference over a pre-registered space of randomization allocations.
@@
We call out the Evaluation & Adaptation step of IGR as an especially critical component of the framework. In some sense, the inclusion of the Evaluation & Adaptation step is a recognition that the study analyst is not all-knowing and hence may not select a good fitness function and restriction rule in their first attempt. In this work, we highlighted three issues to consider in Evaluation & Adaptation: desiderata trade-offs, discriminatory power, and over-restriction. For example, in Vignette 2, the analyst may initially choose to assign higher weight to the interference metric than to the balance metric in order to prioritize reducing interference bias (Figure 4.2.2). After observing the costs that this choice of prioritization incurs in balance, however, the analyst may choose to modify the weighting scheme to instead apply equal weights to both metrics. If the analyst notices that the chosen fitness function produces homogeneous scores or that the restricted assignment mechanism accepts allocations that are highly correlated, then they can make informed updates to the fitness function and to the restriction rule. These iterative refinements push the design towards one that is better and more likely to yield credible inference. As mentioned in Related Works (Section 2), an emphasis on detecting and correcting flawed restricted assignment mechanisms is missing from the standard re-randomization framework.
@@
There are several interesting directions for future work on IGR. Since the fitness function in IGR is hand-crafted, it may be useful to develop theory on how similar the specified fitness function needs to be to the true conditional mean-squared error in order for IGR to produce better effect estimates than naive randomization strategies. When the fitness function is not a faithful proxy, then analysts may choose to use a non-IGR design. Alternatively, analysts could choose to improve the fitness function by using additional sources of data and domain knowledge. If experiments consist of multiple waves or are preceded by a pilot study, then there is opportunity to use information collected from the prior wave or from the pilot study to refine the fitness function used in the next wave or in the actual trial. In Vignette 2, for example, researchers could use a pilot study to estimate the interference network structure within a subsample of the study participants, then extrapolate this network to the entirety of the study sample. Fitness functions for new experiments could potentially also leverage data from prior historical experiments that evaluated different but comparable interventions in similar settings. Lastly, to make IGR as accessible as possible by study analysts, good software for implementation will be useful; this will likely include a dashboard interface that allows users to easily perform each of the steps of IGR, including interacting with visualizations.
Appendix A Simulation Pipeline
@@
    Generate baseline covariates according to a vignette-specific data-generating process.
@@
      For IGR: Enumerate a pool of $M$ candidate allocations and apply a thresholding restriction rule where the threshold is defined as the $m^{th}$ smallest score. Specifically, keep the top $\frac{m}{k}$ scoring allocations and their corresponding $\frac{(k-1)\cdot m}{k}$ mirror allocations (regardless of score) as acceptable allocations for the experiment.
@@
Appendix B Vignette 1 Data-generating Process
We simulate the study data set to encode theory-based covariance between student characteristics (age, gender, and college major), latent traits (ability and confidence), and course performance (homework score and final exam grade). In particular, course performance is assumed to vary with students’ latent ability and confidence, which in turn depend on age, gender, and major. Ability is modeled to increase with age and is higher for some majors than other majors. In the context of STEM majors where the majority of students are men, we model confidence to be higher among men than women. We also model also higher confidence for certain STEM-related majors.
Let $\mathbf{x}_{i}=\begin{pmatrix}age_{i}&gender_{i}&major_{i}&hw_{i}&exam_{i}\end{pmatrix}^{T}$ denote observed baseline covariates for individual $i$ and $\mathbf{u}_{i}=\begin{pmatrix}ability_{i}&confidence_{i}\end{pmatrix}^{T}$ denote latent, unobserved covariates. To generate $gender_{i}$, we randomly set half of the study sample to be men. Although having a fixed gender composition in the study population is not strictly necessary, it makes it easier implement a group formation strategy that assembles groups with certain gender compositions. In a real-world setting, we could construct a study population with a fixed gender ratio by using an enrollment approach that intentionally targets this ratio. We use the following data-generating process for the remaining covariates:
$age_{i}$ $\sim\text{Unif}(19,25)$
$major_{i}$$\sim\{1,2,3\}\text{ with probability }\{0.5,0.3,0.2\}$
$ability_{i}$ $=(age_{i}-\bar{age})/\sigma_{age}-0.5\cdot\mathbbm{1}\{gender_{i}=1\}+\mathbbm{1}\{major_{i}=1\}+\varepsilon$
$confidence_{i}$ $=\mathbbm{1}\{gender_{i}=1\}+\mathbbm{1}\{major_{i}=2\}+\varepsilon$
$hw_{i}$ $=ability_{i}+\epsilon$
$exam_{i}$ $=ability_{i}+confidence_{i}+\varepsilon$
$\varepsilon$ $\sim N(0,1)$
$Y_{i}^{obs}=\text{exam}_{i}+\mathbf{e}_{c(i)}^{T}\boldsymbol{\tau}z_{i}$ (B.1)
where $\mathbf{e}_{c(i)}\in\mathbb{R}^{l}$ is a vector that has a 1 in the $c(i)^{th}$ entry and a 0 everywhere else, and $\boldsymbol{\tau}\in\mathbb{R}^{l}$ is a vector of treatment effects. Entry $j$ in $\boldsymbol{\tau}$ can be interpreted as the effect of the intervention for the composition $r_{a}^{j}$. We run simulations with groups of composition types $\{c_{a}^{1},c_{a}^{2},c_{a}^{3}\}=\{0.5,0.3,0.7\}$ (that is, 50% male, 30% male, and 70% male), with 20 individuals per group. In the Enumeration step of IGR, we enumerate a pool of size $M=100,000$ using a group formation randomization, then apply a thresholding decision rule to keep the $m=500$ allocations with the best scores.
Appendix C Vignette 2 Data-generating Process
@@
$P(A_{ij}=1)$ =$=\begin{cases}\theta^{w}&s(i)=s(j)\\ \theta^{b1}\cdot\tilde{d}(i,j)^{-\gamma}&s(i)\neq s(j),t(i)=t(j)\\ \theta^{b2}\cdot\tilde{d}(i,j)^{-\gamma}&s(i)\neq s(j),t(i)\neq t(j)\\ \end{cases}$ (C.3)
$\tilde{d}(i,j)$ $=\frac{||\mathbf{l}_{s(i)}-\mathbf{l}_{s(j)}||_{2}^{2}}{\min_{(i,j)\in[N]\times[N],s(i)\neq s(j)}||\mathbf{l}_{s(i)}-\mathbf{l}_{s(j)}||_{2}^{2}}$ (C.4)
@@
$Y_{i}^{obs}=\mathbf{X}\boldsymbol{\beta}+\tau z_{i}+\delta\mathbbm{1}\left\{\sum_{j\in\mathcal{N}(i)}z_{j}>q\cdot|\mathcal{N}(i)|\right\}(1-z_{i})$ (C.5)
Note that, under this model, the total average treatment effect $\tau_{TATE}$ is exactly equal to $\tau$. Furthermore, if we use an experimental design that assigns a fixed number of units $N_{c}$ to control, then the bias of a difference-in-means estimator is
$E[\hat{\tau}-\tau]=-\delta\cdot\frac{1}{N_{c}}\sum_{i=1}^{N}\mathbb{E}\left[\mathbbm{1}\left\{\sum_{j\in\mathcal{N}(i)}z_{j}>q\cdot|\mathcal{N}(i)|\right\}(1-z_{i})\right]$

### v2

@@ 7 Discussion
We introduced Inspection-Guided Randomization as a flexible framework for performing restricted randomization with diverse design desiderata and used illustrative examples to showcase its usefulness in different experimental settings. IGR is distinct from prior restricted randomization frameworks along several fronts. It allows for multiple design desiderata, including and beyond balance, thus expanding the use cases for restricted randomization; we showed this with an example of an experiment with interference in Section 4 and a group formation experiment in Appendix E. IGR also separates the process of enumerating candidate allocations – which can be done with an efficient, directed search – from the act of randomization; we explicitly make enumeration and randomization separate steps in the IGR framework. In addition, IGR has a built-in iterative loop for evaluating and adapting the restricted assignment mechanism; we show this in action in Section 4.2, using the help of visual tools described in Appendix D. Lastly, IGR enforces pre-registration of the accepted allocations, thus promoting transparency and reproducibility; in Appendix G, we show that, without pre-registration, it is easy to inflate the Type I error rate above the nominal level.
@@
The core motivation for IGR, as for any restricted randomization, is to prevent the possibility of choosing a bad randomization allocation that leads to an effect estimate that deviates dramatically from the true value of the estimand. For obviously bad allocations, there is little reason not to discard them; this has been well-understood since the early days of the experimental design literature. In a conversation between Cochran and Fisher (recounted by Rubin and Cochran), Fisher is reported to have said that, confronted with a randomized allocation that “exhibited substantial imbalance on a prognostically important baseline covariate,” he would “of course” rerandomize if the experiment had not started yet (Rubin,, 2008).
@@
Historically, the focus of restricted randomization has been on achieving covariate balance in two-arm experiments that compare a single treatment intervention to a control or status quo. However, some interventions, often ones that are behavioral, need to be evaluated through more complicated restricted designs. Researchers across several disciplines, including medicine, education, psychology, and computer science, have recognized this need and have independently applied principles of the IGR framework to designing experiments, albeit without the formal structure of IGR (Murnane et al.,, 2023; Cho et al.,, 2021, 2022; Kizilcec et al.,, 2022; Zahrt et al.,, 2023). Cho et al., (2021) and Cho et al., (2022), for example, used a form of restricted randomization to achieve balance on prognostic covariates in trials with more than two arms. These existing real-life applications suggest that IGR is conceptually straightforward and intuitive, and they also point to how an explicit IGR framework can help to bring consistency to existing efforts to expand the scope of restricted randomization.
@@
In this work, we showed how the IGR framework can be used to design experiments in two settings where existing restricted randomization frameworks are not immediately applicable. In our first example, we applied IGR to the design of experiments with interference. Experimental designs that seek to mitigate interference often do so through a form of cluster randomization, either over naturally existent clusters or over algorithmically detected ones (Ugander et al.,, 2013; Ugander and Yin,, 2023). Without IGR, the study analyst chooses between randomizing at the more aggregated level to limit interference, thus sacrificing power and covariate balance, or randomizing at the more granular level to increase efficiency, thus permitting more interference bias. With IGR, the analyst can construct fitness functions that combine both interference and balance metrics, thus reducing interference bias while maintaining efficiency. IGR can also be used even if we have limited knowledge of interference structure and need to rely on an approximate measure of interference. While we explored a specific setting where interference occurred via neighborhood exposure in a network of interactions, IGR can easily be applied to other interference structures by modifying the inspection metric used to measure the extent of anticipated interference.
In Appendix E, we further show how IGR can be applied to the design of group formation experiments, where effect sizes may exhibit dependence on group composition. Recent work has highlighted that peer context should be taken into consideration as a driver of treatment effect heterogeneity for behavioral and psychological interventions (Yeager et al.,, 2019; Walton and Yeager,, 2020; Kizilcec et al.,, 2020; Kizilcec and Cohen,, 2017). Understanding how an intervention depends on peer context has important implications for understanding causal mechanisms, for devising policies that target certain contexts, and for designing concurrent interventions that can be used to manipulate contexts.
@@
We call out the Evaluation & Adaptation step of IGR as an especially critical component of the framework. In some sense, the inclusion of the Evaluation & Adaptation step is a recognition that the study analyst is not all-knowing and hence may not select a good fitness function and restriction rule in their first attempt. In this work, we highlighted three issues to consider in Evaluation & Adaptation: desiderata trade-offs, discriminatory power, and over-restriction. We introduced visual tools to help with assessing each of these issues (described in detail in Appendix D and applied in Section 4.2). For example, in our simulated experiment, the analyst may initially choose to assign equal weight to the interference metric and the balance metric (Figure 2). After considering the risk of a false null, however, the analyst may choose to modify the weighting scheme to instead upweight the interference metric. These iterative refinements push the design towards one that is better and more likely to yield credible inference. As mentioned in Related Works (Section 2), an emphasis on detecting and correcting flawed restricted assignment mechanisms is missing from standard restricted randomization frameworks such as re-randomization.
Fundamentally, IGR compels the study analyst to invest time and care in the design phase of the experiment. The Specification and Evaluation & Adaptation steps in particular require considerable thoughtfulness, from specifying the study’s design desiderata, to developing the inspection metrics used to measure desiderata satisfaction, to constructing and evaluating the fitness function and restriction rule that characterize the restricted assignment mechanism. Attentiveness to design can produce large payoffs in the bias and precision of effect estimates. While using a tailored estimator in the analysis phase can sometimes produce similar payoffs, focusing efforts on the design phase has additional merits. For one, even the most clever estimators may not be able to salvage a study that was performed under an extremely poor choice of randomization allocation. In addition, as previously mentioned, the Pre-registration step precludes statistical hacking, as long as the analyst is held accountable to doing randomization inference over a pre-registered space of randomization allocations (see Appendix G for an illustration of Type I error rate inflation when the analyst is not held accountable to a pre-registered design and analysis plan).
@@
There are several interesting directions for future work on IGR. Since the fitness function in IGR is hand-crafted, it may be useful to develop theory on how similar the specified fitness function needs to be to the true conditional mean-squared error in order for IGR to produce better effect estimates than naive randomization strategies. When the fitness function is not a faithful proxy, then analysts may choose to use a non-IGR design. Alternatively, analysts could choose to improve the fitness function by using additional sources of data and domain knowledge. If experiments consist of multiple waves or are preceded by a pilot study, then there is opportunity to use information collected from the prior wave or from the pilot study to refine the fitness function used in the next wave or in the actual trial. In an experiment with anticipated interference, for example, researchers could use a pilot study to estimate the interference network structure within a subsample of the study participants, then extrapolate this network to the entirety of the study sample. Fitness functions for new experiments could potentially also leverage data from prior historical experiments that evaluated different but comparable interventions in similar settings. Lastly, to make IGR as accessible as possible by study analysts, good software for implementation will be useful; this will likely include a dashboard interface that allows users to easily perform each of the steps of IGR, including interacting with visualizations.
Appendix A Applications of Tightening, Re-randomization, and Constrained Randomization Use Balance Criteria
Paper Randomization strategy Criteria
Tukey, (1993) Constrained Balance in treatment group sizes within covariate strata
Raab and Butcher, (2001) Constrained L2 balance
Moulton, (2004) Constrained Absolute difference in mean covariate value, geographic balance
Morgan and Rubin, (2012) Re-randomization Mahalanobis distance
de Hoop et al., (2012) Constrained L2 balance
Li et al., (2016) Constrained L2 balance, L1 balance
Branson et al., (2016) Re-randomization Mahalanobis distance
Li et al., (2017) Constrained L2 balance, L1 balance
Basse and Airoldi, (2018) Re-randomization Network balance (e.g. average neighborhood size)
Li et al., (2018) Re-randomization Mahalanobis distance
Ciolino et al., (2019) Constrained Kruskal-Wallis test p-value
Krieger et al., (2020) Re-randomization Absolute difference in mean covariate
Watson et al., (2021) Constrained Trace of the between-arm covariance matrix
Zhou et al., (2021) Constrained Maximum pairwise L2 balance, maximum pairwise Mahalanobis distance
Nordin and Schultzberg, (2022) Re-randomization Mahalanobis distance
Kapelner et al., (2021) Re-randomization Mahalanobis distance
Table A.1: Examples of re-randomization and constrained randomization in the literature. Re-randomization refers to the strategy where allocations are sampled one at a time until the first allocation that satisfies the criterion is drawn. Constrained randomization refers to the strategy where a set of candidate allocations are sampled simultaneously, then filtered down to a set of accepted allocations from which the final allocation is drawn. All of these works use a type of balance criterion (including Basse and Airoldi, 2018 who use a balance criterion based on network properties in order to account for network-correlated outcomes).
Appendix B General Simulation Pipeline
@@
    Generate baseline covariates.
@@
      For IGR: Enumerate a pool of $M$ candidate allocations and apply a thresholding restriction rule where the threshold is defined as the $m^{th}$ smallest score. Specifically, keep the top $\frac{m}{k}$ scoring allocations and possibly their corresponding $\frac{(k-1)\cdot m}{k}$ mirror allocations (regardless of score) as acceptable allocations for the experiment.
@@
Appendix C Data Generation for Experiment With Interference
@@
[…]
@@
$Y_{i}^{obs}=\mathbf{X}\boldsymbol{\beta}+\tau z_{i}+\delta\mathbf{1}\left\{\sum_{j\in\mathcal{N}(i)}z_{j}>q\cdot|\mathcal{N}(i)|\right\}(1-z_{i})$ (C.5)
Note that, under this model, the global average treatment effect $\tau_{TATE}$ is exactly equal to $\tau$. Furthermore, if we use an experimental design that assigns a fixed number of units $N_{c}$ to control, then the bias of a difference-in-means estimator is
$E[\hat{\tau}-\tau]=-\delta\cdot\frac{1}{N_{c}}\sum_{i=1}^{N}\mathbb{E}\left[\mathbf{1}\left\{\sum_{j\in\mathcal{N}(i)}z_{j}>q\cdot|\mathcal{N}(i)|\right\}(1-z_{i})\right]$